7. All of the following problems involve interest that is being compounded continuously (a) Compute the interest rate needed in order to double an invest- ment every seven years. (b) With a 10% interest, how long will it take to triple my initial investment. (c) Irvin forgot what is the interest rate at his bank. All he knows is that after five years, his initial investment doubled. When will it triple? (Hint: There are two unknowns; one is the interest rate Whenever there are two unknowns, two equations are needed. So, use the information about money at different times in an appropriate manner.) (d) A person has money withdrawn from his savings account to be placed in his checking account at a continuous basis. This is at a rate of $10,000 a year. This person started with $20,000 placed in the savings account that has a 5% interest. How much will be in the savings account in one year? (Hint: A different expression for the growth of money needs to be derived. So, start with AM and instead of the Eq. 2.4 expression, determine what ΔΜ equals in this setting. Then, solve the equation.)
Answer #1

a) Interest rate needed to double an investment every seven years

For knowing doubling impact, we should use the 72 principle, formula for computation of rate for an investment to double                         = 72 / Number of years = 72 / 7 = 10.29%

b) With 10% , how long it will take to triple investment

In this case of tripling concept, we should use 114, Number of years = 114 /10 = 11.4 Years

c) Interest rate for doubling amount in 5 years

For knowing doubling impact, we should use the 72 principle, formula for computation of rate for an investment to double                         = 72 / Number of years = 72 / 5 =14.4 %

For knowing within how many years it will triple, we have to used 114 principle = 114 / 14.4 = 7.92 Years

d) Δ M (Savings account balance in the end of 1 year) = Amount in savings account + Interest on savings account – Withdrawable amount every year (10,000)

= $20,000 + 20,000*5% - 10,000 = $11,000

I hope this will help you. If requires clarification, you may comment below.